# Local Thermal Operations and Classical Communication

### v2

@@ Abstract
In quantum thermodynamics, understanding the interplay between locality, thermal constraints, and communication remains an open challenge. In this manuscript, we introduce Local Thermal Operations and Classical Communication (LTOCC), a novel operational framework that unifies the distant laboratories paradigm with thermodynamic restrictions, defining the fundamental limits on transformations between spatially separated systems. We establish a hierarchy of LTOCC protocols, demonstrating inclusion relations between different levels and revealing their deep connection to semilocal thermal operations. To formalize this framework, we develop thermal tensors and bithermal tensors, extending tristochastic tensors to thermodynamic settings and providing new mathematical tools for constrained quantum processes. Furthermore, we explore the role of LTOCC in approximating logical gates such as CNOT and SWAP and investigate its potential to generate correlations between distant systems.
: New J. Phys.
1 Introduction
Amongst the pillars on which modern physics stands, quantum mechanics and thermodynamics stand as two of the most robust ones. As Sir Arthur Eddington has famously stated, “(…) if your theory is found to be against the second law of thermodynamics, I can give you no hope; there is nothing for it but to collapse in deepest humiliation [1].” Similarly, despite significant controversies regarding its ontological status, quantum mechanics has withstood all experimental inquiry it has been subject to over the century of its progress.
Recent years have seen a growth of the efforts focused on joining these two fields in what is called quantum thermodynamics [2, 3, 4], with results applicable to the problems of quantum battery charging [5, 6, 7, 8, 9, 10, 11, 12, 13], cooling [14, 15, 16, 17, 18, 19, 20], autonomous quantum machines [21, 22, 23, 24] and more. Most important in the current context, however, are efforts to provide ultimate bounds on transformations of low-dimensional systems in contact with a heat bath that would respect the conservation of thermal equilibrium. This approach has resulted in the introduction of the resource-theoretic framework of thermal operations, with a wide body of related works already present within the literature [25, 26, 27, 28, 29, 30, 31, 32], extended by a concept of semilocal thermal operations, where local thermal baths in different temperatures interact only via local proxy systems, but interactions between the said systems can be arbitrarily nonlocal, classical or quantum, within the thermodynamical limits [33, 34].
@@ 1 Introduction
In this work, following Sir Eddington’s advice, we propose a framework for operations limited both by locality and communication restricted to classical means, as well as incorporating thermodynamic laws locally for all involved parties, thus adhering to the basic requirements of thermodynamics. For simplicity, we will refer to such operations as local thermal operations and classical communication (LTOCC). In Section 2 we recall all the notions necessary to introduce our framework, including: noisy and thermal operations and cooling maps, local operations and classical communication, and concepts connected to stochastic tensors and multistochasticity. Furthermore, in Section 3 we discuss the concept of semilocal thermal operations (SLTO) and present new results, which are required to establish a connection between SLTO and LTOCC. In Section 4 we introduce the proper framework of LTOCC together with a hierarchy of protocols requiring different amounts of communication and control, demonstrating their relationship with semilocal thermal operations. In Section 5 we study the basic building blocks of the framework – thermal tensors – with a special focus on the nontrivial, symmetric subset of bithermal tensors, which are thermal counterparts of tristochastic tensors. Finally, in Section 6 we present thermal counterparts of standard logical operations which fall within the introduced framework and discuss the correlating abilities of one-round LTOCC and symmetric LTOCC. Section 7 serves as a brief summary of the proposed framework and sets up a list of open problems for future investigation. A and B present a discussion of cooling maps and their alternative derivation, C provides a detailed discussion of one- and two-round LTOCC protocols with and without memory, and D presents new, auxiliary results on semilocal thermal operations and proof of inclusion relation between LTOCC and SLTO.
@@
    Proofs of convexity and closedness under composition for semilocal thermal operations (SLTO) (Theorems 4 and 5)
@@
    Introduction of a new paradigm of local thermal operations with classical communication (LTOCC) and corresponding communication complexity hierarchy, together with proofs of inclusion between different strata of hierarchy (Section 4.3)
@@
    Proof of inclusion of LTOCC with shared randomness in the SLTO in a general quantum scenario for any number of rounds (Theorem 7) and conjecture concerning equality of the sets in the infinite-round limit for energy-incoherent states (Conjecture 1)
@@
    Detailed study of basic building blocks of the framework – thermal and bithermal tensors – with particular attention to the latter as analogues of tristochastic tensors, with preliminary results on the structure of thermal analogue of tristochastic Birkhoff polytope (Section 5)
@@
    Case studies of the use of no-memory LTOCC to approximate CNOT and SWAP logic gates (Section 6.1) and the possibility of creating significant correlations between the involved systems when using memory-extended LTOCC (Section 6.2)
Additionally, in C we present a detailed consideration of 1- and 2-round LTOCC together with placing them in the context of sets of free states for composite resource theories as defined in [39].
2 Preliminaries
2.1 Noisy operations, thermal operations and cooling maps
Let us begin by introducing the essential frameworks that describe the evolution of finite-dimensional quantum states under thermodynamical constraints resulting from contact with a thermal bath, characterised by inverse temperature $\beta=(k_{B}T)^{-1}$ which will also govern local evolution of our subsystem. The primary system of dimension $d$, prepared in an initial state $\rho$ is characterised by its Hamiltonian $H=\sum_{i}E_{i}\outerproduct{E_{i}}{E_{i}}$ |, where we assume the energy levels $E_{i}$ to be non-degenerate. The total system-bath initial state is assumed to be uncorrelated, $\rho\otimes\gamma_{E}$, with the bath characterised by the Hamiltonian $H_{E}$ starting in the Gibbs state defined as
@@ 2.1 Noisy operations, thermal operations and cooling maps
[…]
@@
[$\commutator{U}{H\otimes\mathbbm{1}_{E}+\mathbbm{1}\otimes H_{E}}=0,$ , (3)
@@
[…]
@@
In the following paragraphs, we discuss the resulting restrictions on possible transitions between energy-incoherent states, $\mathbf{p}\rightarrow\mathbf{q}$. Note that the following considerations apply as necesery conditions also whenever the original state $\rho$ is coherent and $\operatorname{diag}(\rho)=\mathbf{p}$ [27].
@@
The finite temperature case of $\beta>0$ has been considered in detail in [27], and here we present only the resulting relation of thermomajorisation. In order to do so, we first begin by defining slope vector $\mathbf{r}(\mathbf{p})$ with entries $r_{i}(\mathbf{p})=p_{i}/\gamma_{i}$. Then, the $\beta$-ordering of a state.
@@
Consider a $d$-dimensional population vector $\mathbf{p}$ and a corresponding Gibbs state $\boldsymbol{\gamma}$. A permutation $\pi_{\mathbf{p}}\in\mathcal{S}$ is called the $\beta$-order of the state $\mathbf{p}$ if it orders the corresponding slope vector in a non-increasing order,
$i<j\Longrightarrow r_{\pi_{\mathbf{p}}^{-1}(i)}(\mathbf{p})\geq r_{\pi_{\mathbf{p}}^{-1}(j)}(\mathbf{p}).$ (6)
Furthermore, we will call the reordered version $p^{\beta}_{i}=p_{\pi(i)}$ as the $\beta$-ordered state.
@@
$\left(\sum_{i=1}^{k}\gamma^{\,\beta}_{i},\leavevmode\nobreak\ \sum_{i=1}^{k}p^{\,\beta}_{i}\right):=\left(\sum_{i=1}^{k}\gamma_{\mathbf{\pi}^{-1}_{\mathbf{p}}(i)},\leavevmode\nobreak\ \sum_{i=1}^{k}p_{\mathbf{\pi}^{-1}_{\mathbf{p}}(i)}\right),$ (7)
@@
$\mathbf{p}\succ_{\beta}\mathbf{q}\iff\forall x\in[0,1]:\leavevmode\nobreak\ f^{\,\beta}_{\mathbf{p}}(x)\geq f^{\,\beta}_{\mathbf{q}}(x)\,.$ (8)
@@
Bistochastic matrices correspond to thermal operations on energy-incoherent states in infinite temperature. However one can also study the opposite zero temperature limit. In the A we summarize the concept of cooling maps, originating for such a study [45], in a scenario with diagonal quantum states. Furthermore in B we show an alternative derivation of cooling maps for diagonal quantum states as a zero temperature limit of Gibbs-preserving matrices.
@@
2.2 Local operations and classical communication
@@
2.3 Multistochastic tensors and multi stochastic Birkhoff polytope
@@
$T(\mathbf{p},\mathbf{q})_{i}=\sum_{jk}T_{ijk}p_{j}q_{k}\leavevmode\nobreak\ ,$ (9)
and the concept can be extended for more distributions by considering tensors of higher rank. To ensure, that the operation is well-defined, all elements of the stochastic tensor must be positive, $T_{ijk}\geq 0$, and sums of its hypercolumns are normalised, $\sum_{i}T_{ijk}=1$, for any $j,k$. Typically, physically interesting phenomena are described using restricted scenarios such as bistochasticity, which appears naturally in noisy operations or as transition probabilities arising from unitary evolution of quantum systems [51, 52]. In the case of tensors instead of one additional condition for the normalisation of rows (hypercolumns in relation to a selected index), one can demand similar conditions to be satisfied by all hypercolumns, which gives rise to multistochastic tensors.
@@
$\forall\;\mathbf{p}\leavevmode\nobreak\ \leavevmode\nobreak\ T(\mathbf{p},\boldsymbol{\eta})=T(\boldsymbol{\eta},\mathbf{p})=\boldsymbol{\eta}$ (10)
@@
For example, sets of both bistochastic matrices and tristochast tensors are convex, forming the so-called Birkhoff polytopes [54]. All extremal bistochastic matrices, vertices of Birkhoff polytope, are permutation matrices [55] – all its entries are either one or zero. Thus one would expect, that permutation tensors, in which all entries are equal to either one or zero, also fully describe the extremal tensors. However, as proven in [56], there are also other extremal multistochastic tensors that do not correspond to permutation tensors. Note, that while layers of permutation tensors are extremal bistochastic matrices, it no longer holds for other extremal multistochastic tensors.
@@
3 Interlude: Auxiliary results on semilocal thermal operations
@@
In our work we treat SLTO as an important background, which operates within minimal assumptions similar to thermal operations, but does not provide any additional structural restrictions on the manner in which the involved systems interact. The exact definition of semilocal thermal operation is as follows:
@@
$\Lambda^{(AB)}(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U(\gamma^{\mathcal{B}^{(A)}}\otimes\gamma^{\mathcal{B}^{(A)}}\otimes\rho^{(AB)})U^{\dagger}\right]$ (13)
@@
$[U,\beta^{(A)}(H^{(A)}+H^{\mathcal{B}^{(A)}})+\beta^{(B)}(H^{(B)}+H^{\mathcal{B}^{(B)}})]=0$ (14b)
@@
The transition between two energy-incoherent states $\rho^{(AB)}\to\sigma^{(AB)}$ can occur under semi-local thermal operation if, and only if, the diagonal of $\rho^{(AB)}$ thermo-majorizes the diagonal of $\sigma^{(AB)}$ with respect to state $\gamma^{(A)}\otimes\gamma^{(B)}$
The above theorem can be extended by an important corollary. Let us perform an atrificial rescaling of Hamiltonian $\tilde{H}^{(A)}=\sqrt{\frac{\beta^{(A)}}{\beta^{(B)}}}H^{(B)},\,\tilde{H}^{(B)}=\sqrt{\frac{\beta^{(B)}}{\beta^{(A)}}}H^{(B)}$, so that the Gibbs states with common inverse temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$ and Hamiltonians $\tilde{H}^{(A)}$ and $\tilde{H}^{(B)}$ are the same as the original Gibbs states $\gamma^{(A)}$ and $\gamma^{(B)}$. Then the condition from Theorem 3, by the Theorem 2, is equivalent to the preservation of global Gibbs state $\gamma^{(AB)}=\gamma^{(A)}\otimes\gamma^{(B)}$ at a common temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$. Thus, we have.
@@
We present two more properties of SLTO, which are crucial from the perspective of resource theories and important in the future work but, to our knowledge, were not proven before as well.
@@
Proofs of Theorems 4 and 5 are provided in D. ∎
@@
4 Framework of Local Thermal Operations and Classical Communication

### v4

@@ Abstract
In quantum thermodynamics, understanding the interplay between locality, thermal constraints, and communication remains an open challenge. In this manuscript, we introduce Local Thermal Operations and Classical Communication (LTOCC), a novel operational framework that unifies the distant laboratories paradigm with thermodynamic restrictions, defining the fundamental limits on transformations between spatially separated systems. We establish a hierarchy of LTOCC protocols, demonstrating inclusion relations between different levels and revealing their deep connection to semilocal thermal operations. To formalize this framework, we develop thermal tensors and bithermal tensors, extending stochastic and tristochastic tensors to thermodynamic settings and providing new mathematical tools for constrained quantum processes. Finally, we present limitations imposed by LTOCC on single- and multi-copy CHSH scenario, demonstrating no violation in former and a gap between thermal and athermal local operations in the latter with respect to their capability to detect entanglement.
Suggested keywords
I Introduction
Amongst the pillars on which modern physics stands, quantum mechanics and thermodynamics present themselves as two of the most robust ones. As Sir Arthur Eddington has famously stated, “(…) if your theory is found to be against the second law of thermodynamics, I can give you no hope; there is nothing for it but to collapse in deepest humiliation [1].” Similarly, despite significant controversies regarding its ontological status, quantum mechanics has withstood all experimental inquiry it has been subject to over the century of its progress.
Recent years have seen a growth of the efforts focused on joining these two fields in what is called quantum thermodynamics [2, 3, 4], with results applicable to the problems of quantum battery charging [5, 6, 7, 8, 9, 10, 11, 12, 13], cooling [14, 15, 16, 17, 18, 19, 20], autonomous quantum machines [21, 22, 23, 24] and more [25]. Most important in the current context, however, are efforts to provide ultimate bounds on transformations of low-dimensional systems in contact with a heat bath that would respect the conservation of thermal equilibrium. This approach has resulted in the introduction of the resource-theoretic framework of thermal operations, with a wide body of related works already present within the literature [26, 27, 28, 29, 30, 31, 32, 33], extended by a concept of semilocal thermal operations, where local thermal baths in different temperatures interact only via local proxy systems, but interactions between the said systems can be arbitrarily nonlocal, classical or quantum, within the thermodynamical limits [34, 35].
@@ I Introduction
In this work, following Sir Eddington’s advice, we propose a framework for operations limited both by locality and communication restricted to classical means, as well as incorporating thermodynamic laws locally for all involved parties, thus adhering to the basic requirements of thermodynamics. For simplicity, we will refer to such operations as local thermal operations and classical communication (LTOCC). In Section II we recall all the notions necessary to introduce our framework, including: noisy and thermal operations and cooling maps, local operations and classical communication, and concepts connected to stochastic tensors and multistochasticity. Furthermore, in Section III we discuss the concept of semilocal thermal operations (SLTO) and present new results, which are required to establish a connection between SLTO and LTOCC. In Section IV we introduce the proper framework of LTOCC together with a hierarchy of protocols requiring different amounts of communication and control, demonstrating their relationship with semilocal thermal operations. In Section V we study the basic building blocks of the framework – thermal tensors – with a special focus on the nontrivial, symmetric subset of bithermal tensors, which are thermal counterparts of tristochastic tensors. Section VI is devoted to study of nonlocal correlations under LTOCC. More specifically, in Section VI.1 we discuss correlation generation under LTOCC from energy-incoherent product states, highlighting significant correlating power of LTOCC with memory. Finally, in Section VI.2 we consider CHSH scenario under LTOCC and demonstrate gap between full and thermally restricted LOCC, thus showing possiblity of detection of athermality usage in the context of Bell nonlocality. Section VII serves as a brief summary of the proposed framework and sets up a list of open problems for future investigation.
Appendix A and Appendix B present a discussion of cooling maps and their alternative derivation, Appendix C provides a detailed discussion of one- and two-round LTOCC protocols with and without memory, and Appendix D presents new, auxiliary results on semilocal thermal operations and proof of inclusion relation between LTOCC and SLTO. Appendix E presents additional auxiliary results on coherence evolution under SLTO and LTOCC protocols. Finally Appendix F discusses realisability of exemplary classical gates, CNOT and SWAP, within the framework of LTOCC.
@@
    Proofs of convexity and closedness under composition for semilocal thermal operations (SLTO) (Theorems 4 and 5) and additional auxiliary results on coherence evolution under SLTO (Appendix E),
@@
    Introduction of a new paradigm of local thermal operations with classical communication (LTOCC) and corresponding communication complexity hierarchy, together with proofs of inclusion between different strata of hierarchy (Section IV.3),
@@
    Proof of inclusion of LTOCC with shared randomness in the SLTO in a general quantum scenario for any number of rounds (Theorem 7) and conjecture concerning equality of the sets in the infinite-round limit for energy-incoherent states (Conjecture 1),
@@
    Detailed study of basic building blocks of the framework – thermal and bithermal tensors – with particular attention to the latter as analogues of tristochastic tensors, with preliminary results on the structure of thermal analogue of tristochastic Birkhoff polytope (Section V),
@@
    Degree of correlations that can be generated under LTOCC with and without memory (Section VI.1)
  6. 6.
    Limitations on Bell nonlocality with LTOCC-restricted operations, including no CHSH inequality breaking in single copy scenario and appearance of gap between LTOCC and Tsirelson bound in multicopy setting (Section VI.2)
Additionally, in Appendix C we present a detailed consideration of 1- and 2-round LTOCC together with placing them in the context of sets of free states for composite resource theories as defined in [40], and in Appendix F we present the limitations of approximating classical logic gates as CNOT and SWAP by LTOCC.
II Preliminaries
II.1 Noisy operations, thermal operations and cooling maps
Let us begin by introducing the essential frameworks that describe the evolution of finite-dimensional quantum states under thermodynamical constraints resulting from contact with a thermal bath, characterised by inverse temperature $\beta=(k_{B}T)^{-1}$ which will also govern local evolution of our subsystem. The primary system of dimension $d$, prepared in an initial state $\rho$ is characterised by its Hamiltonian $|H=\sum_{i}E_{i}\outerproduct{E_{i}}{E_{i}}$, where we assume the energy levels $E_{i}$ to be non-degenerate. The total system-bath initial state is assumed to be uncorrelated, $\rho\otimes\gamma_{E}$, with the bath characterised by the Hamiltonian $H_{E}$ starting in the Gibbs state defined as
@@ II.1 Noisy operations, thermal operations and cooling maps
Hereafter, subscript $E$ denotes the thermal environment. If not stated otherwise, in what follows we will be restricting our considerations of states under thermodynamic evolution to the energy-incoherent states, ie. such that they commute with the Hamiltonian of the primary system, $\commutator{\rho}{H}=0$. With this restriction, there is a one-to-one correspondence between states $|\rho=\sum_{i}p_{i}\outerproduct{E_{i}}{E_{i}}$ and their population vectors $\mathbf{p}=\text{diag}_{H}(\rho)$. Furthermore, we define the thermal equilibrium state of the system $\gamma$} by Eq. (1) with $H_{E}$ replaced with $H$.
@@
$\commutator{U}{H\otimes\mathbbm{1}_{E}+\mathbbm{1}\otimes H_{E}}=0,$ (3)
@@
In order to provide proper background to the main topic of this manuscript we need to consider the conditions for the convertibility of states. In the case of finite temperatures, given a pair of states $\rho$ and $\sigma$, the transition is possible if and only if there exists a thermal operation $\mathcal{E}$ such that $\mathcal{E}(\rho)=\sigma$ and $\mathcal{E}(\gamma)=\gamma$. However, as we limit ourselves to energy-incoherent states, $\commutator{\rho}{H}=\commutator{\sigma}{H}=0$, the condition reduces to existence of a stochastic matrix $\Lambda$ acting on the respective populations $\mathbf{p}=\operatorname{diag}_{H}(\rho),\,\mathbf{q}=\operatorname{diag}_{H}(\sigma)$ such that [27]
@@
In the following paragraphs, we discuss the resulting restrictions on possible transitions between energy-incoherent states, $\mathbf{p}\rightarrow\mathbf{q}$. Note that the following considerations apply as necessary conditions also whenever the original state $\rho$ is coherent and $\operatorname{diag}(\rho)=\mathbf{p}$ [28].
@@
The finite temperature case of $\beta>0$ has been considered in detail in [28], and here we present only the resulting relation of thermomajorisation. together with the corresponding notion of $\beta$-ordering.
@@
Consider a $d$-dimensional population vector $\mathbf{p}$ and a corresponding Gibbs state $\boldsymbol{\gamma}$. A permutation $\pi_{\mathbf{p}}\in\mathcal{S}$ is called a $\beta$-order of the state $\mathbf{p}$ if it orders the ratios between populations of $\mathbf{p}$ and $\boldsymbol{\gamma}$ in a non-increasing order,
$i<j\Longrightarrow\frac{p_{\pi_{\mathbf{p}}(i)}}{\gamma_{\pi_{\mathbf{p}}(i)}}\geq\frac{p_{\pi_{\mathbf{p}}(j)}}{\gamma_{\pi_{\mathbf{p}}(j)}}.$ (6)
Furthermore, we will call the reordered version $p^{\beta}_{i}=p_{\pi_{\mathbf{p}}}(i)$ as the $\beta$-ordered state.
@@
$\left(\sum_{i=1}^{k}\gamma^{\,\beta}_{i},\penalty 10000\ \sum_{i=1}^{k}p^{\,\beta}_{i}\right):=\left(\sum_{i=1}^{k}\gamma_{\mathbf{\pi}_{\mathbf{p}}(i)},\penalty 10000\ \sum_{i=1}^{k}p_{\mathbf{\pi}_{\mathbf{p}}(i)}\right),$ (7)
@@
$\mathbf{p}\succ_{\beta}\mathbf{q}\iff\forall x\in[0,1]:\penalty 10000\ f^{\,\beta}_{\mathbf{p}}(x)\geq f^{\,\beta}_{\mathbf{q}}(x)\,.$ (8)
@@
Bistochastic matrices correspond to thermal operations on energy-incoherent states in infinite temperature. However one can also study the opposite zero temperature limit. In Appendix A we summarize the concept of cooling maps, originating for such a study [46], in a scenario with diagonal quantum states. Furthermore in Appendix B we show an alternative derivation of cooling maps for diagonal quantum states as a zero temperature limit of Gibbs-preserving matrices.
@@
II.2 Local operations and classical communication
@@
II.3 Multistochastic tensors and multi stochastic Birkhoff polytope
@@
$T(\mathbf{p},\mathbf{q})_{i}=\sum_{jk}T_{ijk}p_{j}q_{k}\penalty 10000\ ,$ (9)
and the concept can be extended for more distributions by considering tensors of higher rank. To ensure, that the operation is well-defined, all elements of the stochastic tensor must be positive, $T_{ijk}\geq 0$, and sums of its hypercolumns are normalised, $\sum_{i}T_{ijk}=1$, for any $j,k$.
Stochastic tensors arise naturally in the study of higher-order Markov chains [52, 53, 54], in which the next probability distribution $\mathbf{p}^{(n+1)}$ depends not only on the current $\mathbf{p}^{(n)}$, but also previous one $\mathbf{p}^{(n-1)}$:
$p_{i}^{(n+1)}=\sum_{j,k}T_{ijk}p_{j}^{(n)}p_{k}^{(n-1)}.$
Typically, physically interesting phenomena are described using restricted scenarios such as bistochasticity, which appears naturally in noisy operations or as transition probabilities arising from unitary evolution of quantum systems [55, 56]. In the case of tensors instead of one additional condition for the normalisation of rows (hypercolumns in relation to a selected index), one can demand similar conditions to be satisfied by all hypercolumns, which gives rise to multistochastic tensors.
@@
$\forall\;\mathbf{p}\penalty 10000\ \penalty 10000\ T(\mathbf{p},\boldsymbol{\eta})=T(\boldsymbol{\eta},\mathbf{p})=\boldsymbol{\eta}$ (10)
These extra properties were leveraged in works such as [57, 58] to generalize the notion of convolution of probability vectors
$\mathbf{p}*\mathbf{q}=T(p,q),$
with $T_{ijk}=\delta_{j+k-i}$ in stardard case, and subsequently obtain its quantum analogues.
@@
For example, sets of both bistochastic matrices and tristochastic tensors are convex, forming the so-called Birkhoff polytopes [60]. All extremal bistochastic matrices, vertices of Birkhoff polytope, are permutation matrices [61] – all its entries are either one or zero. Thus one would expect, that permutation tensors, in which all entries are equal to either one or zero, also fully describe the extremal tensors. However, as proven in [62], there are also other extremal multistochastic tensors that do not correspond to permutation tensors. Note, that while layers of permutation tensors are extremal bistochastic matrices, it no longer holds for other extremal multistochastic tensors.
@@
III Interlude: Auxiliary results on semilocal thermal operations
@@
In our work we treat SLTO as an important background, which operates within minimal assumptions similar to thermal operations involving two systems in contact with their own local thermal environments, but does not provide any additional structural restrictions on the manner in which the involved systems interact. The exact definition of semilocal thermal operation is as follows:
@@
$\Lambda^{(AB)}\quantity(\rho^{(AB)})=\text{Tr}_{\mathcal{B}^{(A)},\mathcal{B}^{(B)}}\left[U(\gamma^{\mathcal{B}^{(A)}}\otimes\gamma^{\mathcal{B}^{(B)}}\otimes\rho^{(AB)})U^{\dagger}\right]$ (13)
@@
$\quantity[U,\beta^{(A)}(H^{(A)}+H^{\mathcal{B}^{(A)}})+\beta^{(B)}(H^{(B)}+H^{\mathcal{B}^{(B)}})]=0$ (14b)
@@
The transition between two energy-incoherent states $\rho^{(AB)}\to\sigma^{(AB)}$ can occur under semi-local thermal operation if, and only if, the diagonal of $\rho^{(AB)}$ thermo-majorizes the diagonal of $\sigma^{(AB)}$ with respect to state $\gamma^{(A)}\otimes\gamma^{(B)}$.
The above theorem can be extended by an important corollary. Let us perform an artificial rescaling of Hamiltonian $\tilde{H}^{(A)}=\sqrt{\frac{\beta^{(A)}}{\beta^{(B)}}}H^{(B)},\,\tilde{H}^{(B)}=\sqrt{\frac{\beta^{(B)}}{\beta^{(A)}}}H^{(B)}$, so that the Gibbs states with common inverse temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$ and Hamiltonians $\tilde{H}^{(A)}$ and $\tilde{H}^{(B)}$ are the same as the original Gibbs states $\gamma^{(A)}$ and $\gamma^{(B)}$. Then the condition from Theorem 3, by the Theorem 2, is equivalent to the preservation of global Gibbs state $\gamma^{(AB)}=\gamma^{(A)}\otimes\gamma^{(B)}$ at a common temperature $\tilde{\beta}=\sqrt{\beta^{(A)}\beta^{(B)}}$. Thus, we have.
@@
The study of SLTO action on coherent states remains largely unexplored, thus in Appendix E we develop preliminary result in this direction, which will prove useful in a later part of the manuscript.
Finally, we present two additional properties of SLTO, which are crucial from the perspective of resource theories and important in further part of this work but, to our knowledge, were not proven before.
@@
Proofs of Theorems 4 and 5 are provided in Appendix D. ∎
@@
IV Framework of Local Thermal Operations and Classical Communication